OpenStudy (anonymous):

What is the value of the summation \[\Sigma i =14(2i +6i2)\] ? A)180 B)200 C)220 D)240

OpenStudy (anonymous):

@phi

OpenStudy (phi):

can you use the equation editor?

OpenStudy (anonymous):

that didnt come out right

OpenStudy (anonymous):

\[\Sigma i =14(2i +6i2)\]

OpenStudy (anonymous):

that is what it's supposed to be

OpenStudy (phi):

that does not make sense. summations have a start and end value for example \[\sum_{i=1}^{3}i= 1+2+3\]

OpenStudy (anonymous):

\[\sum_{i=1}^{4} (2i+6i^{2})\]

OpenStudy (anonymous):

thats the real equation

OpenStudy (phi):

ok , write it like this \[2\sum_{i=1}^{4} i+6\sum_{i=1}^{4} i^{2} \] there are "closed-form" solutions to these summations https://en.wikipedia.org/wiki/Triangular_number https://en.wikipedia.org/wiki/Square_pyramidal_number#Formula

OpenStudy (phi):

or, with only 4 terms, you could do it "by hand"

OpenStudy (phi):

the first summation is the sum of 1 to 4 then multiply that sum by 2 the second summation is the sum of the first four squares 1,4,9,16 I assume you can add those up. then multiply by 6 finally, add the two different numbers for the combined answer

OpenStudy (anonymous):

Thanks so much @phi you da best